# Thermoelectric energy conversion in molecular junctions out of equilibrium

### v1

@@ IV Benchmark simulations for a quantum-dot system
Figure 5: Time-dependent energy current $J_{L}^{E}$ at the left-lead interface with the quantum dot being coupled to the leads with $\gamma=0.5$. (a) High temperature $\mu_{L}=\mu_{R}=0$, $\beta_{L}=\beta_{R}=1$; (b) chemical potential drop $\mu_{L}=-\mu_{R}=0.5,\beta_{L}=\beta_{R}=10$; and (c) temperature gradient $\mu_{L}=-\mu_{R}=0.5,\beta_{L}=1,\beta_{R}=10$. Insets display the equilibrium energy current, before the sudden gate voltage is switched on, in terms of the Lorentzian bandwidth $\Omega$.
@@
With charge densities and currents showing excellent agreement across different levels of approximation and with the earlier literature, we turn to the energy-current response in the intermediate coupling regime $\gamma=0.5$, modifying the transport setup accordingly. Fig. 5 displays three different cases (a) aligned chemical potentials $\mu_{L}=\mu_{R}=0$ at high temperature $\beta_{L}=\beta_{R}=1$; (b) chemical potential drop $\mu_{L}=-\mu_{R}=0.5$ at lower temperature $\beta_{L}=\beta_{R}=10$; and (c) chemical potential drop $\mu_{L}=-\mu_{R}=0.5$ with temperature gradient $\beta_{L}=1,\beta_{R}=10$. In all cases, the pole count is set to $N_{p}=40$. The GKBA (WBLA) result is again shifted the most from the equilibrium value. For GKBA, the finite-width Lorentzians capture the correct trend well in all cases, but there is always a shift in the equilibrium value. It is worth pointing out that now the cases in Fig. 5(b) and 5(c) display chemical potential drops and thermal gradients, i.e., the equilibrium energy current being non-zero is physically expected. In all cases, the $i$GKBA approach corrects for the unphysical behavior observed in GKBA. In the insets of Fig. 5 we also show the dependence on the Lorentzian bandwidth $\Omega$. For the equilibrium energy current, the GKBA results are not only unphysical but the convergence toward the WBLA at $\Omega\to\infty$ is fairly slow. In contrast, the $i$GKBA approach accurately resolves the equilibrium energy current, which does not vary significantly with the bandwidth $\Omega$.
@@
Figure 6: Short-time transient behavior of the energy current. The WBLA result displays an abrupt jump at the voltage switch-on time, which can be continuously modeled by applying wider bandwidth Lorentzians. The broadening parameter $\Omega\in[2,100]$ is shown in varying color, from magenta to cyan.
@@
The broadening parameter $\Omega$ in the finite-width Lorentzian has a pronounced impact on the initial transient dynamics, as already evidenced in Figs. 3-5. To further illustrate this behavior, Fig. 6 zooms into the early-time regime of the energy current shown in Fig. 5(a), immediately following the gate-voltage quench at $t=0$. With the analytic results at the WBLA [81], we see that the energy current exhibits an abrupt jump at the time of the voltage switch-on. The jump can be shown to build up continuously by applying a larger broadening $\Omega$ in the $i$GKBA scheme. Similar findings have also been reported in the case of thermomechanical potentials across the junction [63]. At larger times $t\gtrsim 1$, the energy currents calculated via large-bandwidth Lorentzians again resolve the WBLA result accurately.
@@ V Comparison of ($i$)GKBA with the full Kadanoff-Baym equations
We now consider a molecular junction modeled at the Pariser-Parr-Pople [82, 83] (PPP) level, where the kinetic and interaction matrix elements are semi-empirically obtained by fitting to more sophisticated calculations. Specifically, we study a cyclobutadiene molecule attached to donor-acceptor-like leads [84]. This molecule consists of four atomic sites arranged in a ring (Fig. 1) and is modeled using PPP parameters derived from an effective valence shell Hamiltonian [85]. Referring to Eq. (1), the single-particle matrix (in atomic units) is given by
@@
The slight asymmetry in the hopping and interaction matrix elements arises from the fact that cyclobutadiene forms a rectangular rather than a square structure [86]. We note in passing that the physico-chemical properties of the cyclobutadiene molecule are fairly multifaceted [87]. Our aim here is to assess the validity of the $i$GKBA scheme when compared against the Kadanoff-Baym equations, independent of how accurate the PPP model description is, and we turn aside further discussions on, e.g., the antiaromatic nature and the resulting instability of the molecule [88].
@@
The molecular junction is driven out of equilibrium by a sudden bias voltage on the leads, $V_{\alpha}(t)=V_{\alpha}^{0}\theta(t)$, and we consider weak ($V_{L}^{0}=-V_{R}^{0}=\Delta/8$) and strong ($V_{L}^{0}=-V_{R}^{0}=\Delta/2$) bias voltages. In these situations, the energy centroids of the leads are set to $\epsilon_{L/R}=\mu\pm\Delta/8$ for the weak driving and $\epsilon_{L/R}=\mu\pm\Delta/2$ for the strong driving. The driving parameters are chosen to highlight the capability of the $i$GKBA scheme [61], which does not rely on the uniformity of the lead density of states. The strong driving case is enough to excite a sizable electric current across the HOMO-LUMO gap.
We include the effect of two-body interaction at the Hartree-Fock level, i.e., we neglect the correlation self-energy in Eq. (II.2). This simplified modeling of the molecular junction allows us to explore the new approach with mathematical transparency, facilitating the comparison of the $i$GKBA approach with the full Kadanoff-Baym equations.
@@
In Fig. 7, we present comparative simulations of the various methods for the time-dependent charge current at the right-lead interface. The benchmark KBE results are based on the method of Ref. [41], which we supplement with the Lorentzian line-width function for the embedding self-energies, in order to consistently compare it with the present development. We take the finite-bandwidth Lorentzians as $\Omega=2\Delta$. In contrast to the adiabatic switching procedure in the ($i$)GKBA methods, the KBE approach determines the coupled equilibrium state by solving the Dyson equation on the imaginary time axis [89, 90, 91]. In practice, the imaginary time grid, $\tau\in[0,-\beta]$, is discretized using a uniform power mesh [92, 93] with $p=6$, $u=8$ (total number of $\tau$-points being $12pu+$ 1) to achieve convergence in the total energy up to the relative error $10^{-6}$ for all the cases considered.
In all the cases reported in Fig. 7, we see that $i$GKBA systematically improves upon the GKBA result, and the agreement with KBE is mostly excellent. While the initial transient response to the sudden bias voltage and the following oscillations are qualitatively similar in all methods, quantitative differences appear. The most prominent improvements of the $i$GKBA are the slow decay toward the stationary state and the value of the stationary current. These are both incorrectly described by the GKBA, especially in the strong-bias cases.
@@ VI Time-resolved thermoelectric energy conversion
Now that the validity range of the $i$GKBA approach has been established, we continue with thermoelectric simulations for the molecular junction of the previous section. In addition to a bias voltage, we apply a temperature gradient with $\beta_{L}\neq\beta_{R}$ over the junction. To address thermoelectric energy conversion, we look at the total charge and heat currents $J^{(H)}_{\mathrm{tot}}=(J^{(H)}_{R}-J^{(H)}_{L})/2$ with $J^{H}_{\alpha}=J^{E}_{\alpha}-\mu_{\alpha}J_{\alpha}$. In the linear-response regime, the relationship between the two is commonly written as [7, 1]
@@
where $\mathcal{G}$ and $\mathcal{K}$ are the electrical and thermal conductivities, respectively, and $\mathcal{L}$, $\mathcal{R}$ are the thermoelectric coefficients related to the Seebeck and Peltier effects (connected by the Onsager relations), respectively. Here, $\delta V=V_{L}^{0}-V_{R}^{0}$ and $\delta T=1/\beta_{L}-1/\beta_{R}$. For a more realistic driving protocol, we apply a time-dependent voltage profile, $V_{\alpha}(t)=V_{\alpha}^{0}/[1+\exp(-\kappa t)]$ with $\kappa=5\Delta$, which means that around $t=0$ the bias rapidly yet smoothly increases from 00 to $V_{\alpha}^{0}$.
From Eq. (37) we infer that in the absence of charge current, we may calculate [94]
$S\equiv\frac{\mathcal{L}}{\mathcal{G}}=-\left.\frac{\delta V}{\delta T}\right|_{J_{\mathrm{tot}}=0},$ (38)
which is known as the Seebeck coefficient or thermopower. This enables us to determine time-resolved thermopower by adjusting the bias voltage so that the charge current vanishes. Applying a thermal gradient over the junction, $\delta T>0$, generates a non-zero charge current, which can be ‘countered’ by applying a voltage $\delta V<0$. In practice, this type of thermoelectric energy harvesting involves driving an electric current against a load using a thermal gradient as a driving force [10]. Our motivation here is for dynamically enhanced thermoelectric efficiency in a far-from-equilibrium setting, possibly surpassing steady-state limitations set by static material properties and compatibility conditions [95, 96, 97].
@@
The time-resolved thermopower is shown in Fig. 8(b) for different values of the coupling strength $\gamma^{0}$ and the lead temperatures $(\beta_{L},\beta_{R})$. During the adiabatic switching period ($t<0$), in the absence of applied voltages, the charge currents have saturated to certain non-zero values. The thermopower is then evaluated for times $t>0$ only. It exhibits a rapid drop within a few $\Delta^{-1}$ due to fast charge fluctuations in the molecule. For $\gamma^{0}=\Delta/10$ and $(\beta_{L},\beta_{R})=(8,40)/\Delta$, we see a clear transient enhancement of the thermopower, similar to the one reported in Ref. [94]. The transient behavior thus indicates an increased efficiency of the conversion of thermoelectric energy. Since the parameter space of our thermoelectric transport setup is very large, we do not aim to study exhaustively the optimal conditions for this effect but defer this to future work. In addition, the determined thermopower curves include standard deviations originating from the linear-fit procedure explained above, and we observe that the linear relationship holds very accurately throughout the transient dynamics.
@@
where $J_{L}^{H}$ is the heat current from the left (hot) lead to the molecule. This formulation allows us to evaluate $\eta$ without requiring explicit knowledge of either $R_{s}$ or $R_{l}$. Instead, we parametrize the system in terms of the observable load voltage $\delta V$, which can be held approximately constant even in the transient regime by connecting a capacitor in parallel to the load, and which we scan over the interval $[0,V_{o}]$. Due to the time-dependent nature of the setup, we set $V_{o}$ equal to the mean value of the extracted Seebeck voltages shown in Fig. 8 [78]. This captures both the short- and open-circuit limits, where the efficiency necessarily vanishes due to zero extracted power. The efficiency reaches a maximum at an intermediate value of $\delta V$, allowing a full characterization of the junction performance based solely on measurable quantities.
@@
Overall, the calculated efficiencies are relatively low ($\lesssim 10$%), which is attributed to the specific molecular structure, coupling configuration, and the applied temperature gradient. For voltages outside the range $[0,V_{o}]$, energy is effectively pumped into the system, and the efficiency drops significantly or becomes negative, as dictated by the direction of currents and the form of Eq. (39). More optimized conditions for efficient thermoelectric energy conversion could be explored within the parameter space of this model, or by considering alternative molecular junction setups, and we defer more thorough investigations of quantum thermodynamics phenomena [98, 99] to future work. Ultimately, the Carnot efficiency $\eta_{C}=1-T_{R}/T_{L}$ sets the fundamental upper limit. This bound is only reached in the idealized limit where the thermoelectric figure of merit $ZT=\mathcal{G}S^{2}T/\mathcal{K}\to\infty$; for any finite $ZT$, the efficiency remains strictly below $\eta_{C}$ [100]. Nonetheless, we observe that the transiently enhanced efficiency can clearly surpass the stationary value.
@@ VII Conclusion
Applying the GKBA iteratively in the reconstruction equation resulted in a set of 14 coupled equations of motion, referred to as the $i$GKBA scheme. Starting from the full $i$GKBA scheme, different levels of approximation were explored. Neglecting correction terms in the collision integral reduced to the GKBA framework, still incorporating retarded/advanced propagators beyond the wide-band limit. We exemplified the different levels of approximation by focusing on thermoelectric transport in a two-terminal quantum dot device, for which analytical benchmark results were available. The conventional GKBA proved insufficient for energy current calculations due to an unphysical divergence in the WBLA, which was resolved by the finite-bandwidth $i$GKBA scheme. We also found that the short-time transient behavior of the energy current exhibited an abrupt jump, which was correctly resolved by wider-bandwidth Lorentzians in the $i$GKBA scheme.
Finally, we modeled thermoelectric transport in a cyclobutadiene molecular junction — an interacting system — analyzing the non-equilibrium dynamics under both weak and strong coupling, as well as varying temperatures. By comparing with the numerical solution of the full Kadanoff-Baym equations, the $i$GKBA scheme was found to systematically improve upon GKBA by more accurately capturing relaxation dynamics and steady-state currents. Using the transient charge and heat currents, we investigated thermoelectric energy conversion by extracting the time-resolved Seebeck coefficient (thermopower) and the device efficiency. The molecular junction, when operating in the transient regime, exhibited a clear enhancement in thermoelectric conversion efficiency compared to its stationary-state performance.

### v2

@@ IV Benchmark simulations for a quantum-dot system
Figure 5: Time-dependent energy current $J_{L}^{E}$ at the left-lead interface with the quantum dot being coupled to the leads with $\gamma=\varepsilon/2$, and driven by a sudden gate voltage $u(t)=(5\varepsilon/2)\theta(t)$. (a) High temperature $\mu_{L}=\mu_{R}=0$, $\beta_{L}=\beta_{R}=1/\varepsilon$; (b) chemical potential drop $\mu_{L}=-\mu_{R}=\varepsilon/2,\beta_{L}=\beta_{R}=10/\varepsilon$; and (c) temperature gradient $\mu_{L}=-\mu_{R}=\varepsilon/2,\beta_{L}=1/\varepsilon,\beta_{R}=10/\varepsilon$. Insets display the equilibrium energy current, before the sudden gate voltage is switched on, in terms of the Lorentzian bandwidth $\Omega$.
@@
With charge densities and currents showing excellent agreement across different levels of approximation and with the earlier literature, we turn to the energy-current response in the intermediate coupling regime $\gamma=\varepsilon/2$, modifying the transport setup accordingly. Fig. 5 displays three different cases (a) aligned chemical potentials $\mu_{L}=\mu_{R}=0$ at high temperature $\beta_{L}=\beta_{R}=1/\varepsilon$; (b) chemical potential drop $\mu_{L}=-\mu_{R}=\varepsilon/2$ at lower temperature $\beta_{L}=\beta_{R}=10/\varepsilon$; and (c) chemical potential drop $\mu_{L}=-\mu_{R}=\varepsilon/2$ with temperature gradient $\beta_{L}=1/\varepsilon,\beta_{R}=10/\varepsilon$. In all cases, the pole count is set to $N_{p}=40$. The GKBA (WBLA) result is again shifted the most from the equilibrium value. For GKBA, the finite-width Lorentzians capture the correct trend well in all cases, but there is always a shift in the equilibrium value. It is worth pointing out that now the cases in Fig. 5(b) and 5(c) display chemical potential drops and thermal gradients, i.e., the equilibrium energy current being non-zero is physically expected. In all cases, the iiGKBA approach corrects for the unphysical behavior observed in GKBA. In the insets of Fig. 5 we also show the dependence on the Lorentzian bandwidth $\Omega$. For the equilibrium energy current, the GKBA results are not only unphysical but the convergence toward the WBLA at $\Omega\to\infty$ is fairly slow. In contrast, the iiGKBA approach accurately resolves the equilibrium energy current, which does not vary significantly with the bandwidth $\Omega$.
@@
Figure 6: Short-time transient behavior of the energy current. The WBLA result displays an abrupt jump at the voltage switch-on time, which can be continuously modeled by applying wider bandwidth Lorentzians. The broadening parameter $\Omega/\varepsilon\in[2,100]$ is shown in varying color, from magenta to cyan.
@@
The broadening parameter $\Omega$ in the finite-width Lorentzian has a pronounced impact on the initial transient dynamics, as already evidenced in Figs. 3-5. To further illustrate this behavior, Fig. 6 zooms into the early-time regime of the energy current shown in Fig. 5(a), immediately following the gate-voltage quench at $t=0$. With the analytic results at the WBLA [90], we see that the energy current exhibits an abrupt jump at the time of the voltage switch-on. The jump can be shown to build up continuously by applying a larger broadening $\Omega$ in the iiGKBA scheme. Similar findings have also been reported in the case of thermomechanical potentials across the junction [74]. At larger times $t\gtrsim\varepsilon^{-1}$, the energy currents calculated via large-bandwidth Lorentzians again resolve the WBLA result accurately.
@@ V Comparison of ($i$)GKBA with the full Kadanoff-Baym equations
We now consider an interacting molecular junction modeled at the Pariser-Parr-Pople [91, 92] (PPP) level, where the kinetic and interaction matrix elements are semi-empirically obtained by fitting to more sophisticated calculations. Specifically, we study a cyclobutadiene molecule attached to donor-acceptor-like leads [93]. This molecule consists of four atomic sites arranged in a ring (Fig. 1) and is modeled using PPP parameters derived from an effective valence shell Hamiltonian [94]. Referring to Eq. (1), the single-particle matrix (in atomic units) is given by
@@
The slight asymmetry in the hopping and interaction matrix elements arises from the fact that cyclobutadiene forms a rectangular rather than a square structure [95]. We note in passing that the physico-chemical properties of the cyclobutadiene molecule are fairly multifaceted [96]. Our aim here is to assess the validity of the iiGKBA scheme when compared against the Kadanoff-Baym equations, independent of how accurate the PPP model description is, and we turn aside further discussions on, e.g., the antiaromatic nature and the resulting instability of the molecule [97].
@@
The molecular junction is driven out of equilibrium by a sudden bias voltage on the leads, $V_{\alpha}(t)=V_{\alpha}^{0}\theta(t)$, and we consider weak ($V_{L}^{0}=-V_{R}^{0}=\Delta/8$) and strong ($V_{L}^{0}=-V_{R}^{0}=\Delta/2$) bias voltages. In these situations, the energy centroids of the leads are set to $\epsilon_{L/R}=\mu\pm\Delta/8$ for the weak driving and $\epsilon_{L/R}=\mu\pm\Delta/2$ for the strong driving. The driving parameters are chosen to highlight the capability of the iiGKBA scheme [72], which does not rely on the uniformity of the lead density of states. The strong driving case is enough to excite a sizable electric current across the HOMO-LUMO gap.
We include the effect of two-body interaction at the Hartree-Fock level, i.e., we neglect the correlation self-energy in Eq. (II.2). This simplified modeling of the molecular junction allows us to explore the new approach with mathematical transparency, facilitating the comparison of the iiGKBA approach with the full Kadanoff-Baym equations.
@@
In Fig. 7, we present comparative simulations of the various methods for the time-dependent charge current at the right-lead interface. The benchmark KBE results are based on the method of Ref. [49], which we supplement with the Lorentzian line-width function for the embedding self-energies, in order to consistently compare it with the present development. We take the finite-bandwidth Lorentzians as $\Omega=2\Delta$. In contrast to the adiabatic switching procedure in the ($i$)GKBA methods, the KBE approach determines the coupled equilibrium state by solving the Dyson equation on the imaginary time axis [98, 99, 100]. In practice, the imaginary time grid, $\tau\in[0,-\beta]$, is discretized using a uniform power mesh [101, 102] with $p=6$, $u=8$ (total number of $\tau$-points being $2pu+1$) to achieve convergence in the total energy up to the relative error $10^{-6}$ for all the cases considered.
In all the cases reported in Fig. 7, we see that iiGKBA systematically improves upon the GKBA result, and the agreement with KBE is mostly excellent. While the initial transient response to the sudden bias voltage and the following oscillations are qualitatively similar in all methods, quantitative differences appear. The most prominent improvements of the iiGKBA are the slow decay toward the stationary state and the value of the stationary current. These are both incorrectly described by the GKBA, especially in the strong-bias cases.
@@ VI Time-resolved thermoelectric energy conversion
Now that the validity range of the iiGKBA approach has been established, we continue with thermoelectric simulations for the molecular junction of the previous section. In addition to a bias voltage, we apply a temperature gradient with $\beta_{L}\neq\beta_{R}$ over the junction. To address thermoelectric energy conversion, we look at the total charge and heat currents $J^{(H)}_{\mathrm{tot}}=(J^{(H)}_{R}-J^{(H)}_{L})/2$ with $J^{H}_{\alpha}=J^{E}_{\alpha}-\mu_{\alpha}J_{\alpha}$. In the linear-response regime, the relation between currents and driving fields can be expressed in matrix form as [10, 4]
@@
where $\mathcal{G},\mathcal{L},\mathcal{R},\mathcal{K}$ denote Onsager’s kinetic coefficients. The physical transport coefficients, such as electrical and thermal conductivities, and the Seebeck and Peltier coefficients, are defined in terms of these [103]. From Eq. (37) we obtain, in the absence of charge current [104],
$S\equiv-\left.\frac{\delta V}{\delta T}\right|_{J_{\mathrm{tot}}=0}=\frac{\mathcal{L}}{\mathcal{G}},$ (38)
which defines the Seebeck coefficient, or thermopower. This relation then enables the calculation of time-resolved thermopower by adjusting the bias voltage so that the charge current vanishes, thereby providing a basis for characterizing thermoelectric energy conversion in the molecular junction under non-stationary conditions.
For a more realistic driving protocol, we apply a time-dependent voltage profile, $V_{\alpha}(t)=V_{\alpha}^{0}/[1+\exp(-\kappa t)]$ with $\kappa=5\Delta$, which means that around $t=0$ the bias rapidly yet smoothly increases from 0 to $V_{\alpha}^{0}$. Applying a thermal gradient over the junction, $\delta T\equiv 1/\beta_{L}-1/\beta_{R}>0$, generates a non-zero charge current, which can be ‘countered’ by applying a voltage $\delta V\equiv V_{L}^{0}-V_{R}^{0}<0$. In practice, this type of thermoelectric energy harvesting involves driving an electric current against a load using a thermal gradient as a driving force [14]. Our motivation here is for dynamically enhanced thermoelectric efficiency in a far-from-equilibrium setting, possibly surpassing steady-state limitations set by static material properties and compatibility conditions [105, 106, 107].
@@
The time-resolved thermopower is shown in Fig. 8(b) for different values of the coupling strength $\gamma^{0}$ and the lead temperatures $(\beta_{L},\beta_{R})$. During the adiabatic switching period ($t<0$), in the absence of applied voltages, the charge currents have saturated to certain non-zero values. The thermopower is then evaluated for times $t>0$ only. It exhibits a rapid drop within a few $\Delta^{-1}$ due to fast charge fluctuations in the molecule. For $\gamma^{0}=\Delta/10$ and $(\beta_{L},\beta_{R})=(8,40)/\Delta$, we see a clear transient enhancement of the thermopower, similar to the one reported in Ref. [104]. The transient behavior thus indicates an increased efficiency of the conversion of thermoelectric energy. Since the parameter space of our thermoelectric transport setup is very large, we do not aim to study exhaustively the optimal conditions for this effect but defer this to future work. In addition, the determined thermopower curves include standard deviations originating from the linear-fit procedure explained above, and we observe that the linear relationship holds very accurately throughout the transient dynamics. Similar to the extraction of the time-resolved thermopower, also the thermoelectric figure of merit, $ZT\equiv\mathcal{G}S^{2}T/\mathcal{K}$, may be obtained from a similar analysis based on the charge and heat current data [87].
@@
where $J_{L}^{H}$ is the heat current from the left (hot) lead to the molecule. This formulation allows us to evaluate $\eta$ without requiring explicit knowledge of either $R_{s}$ or $R_{l}$. Instead, we parametrize the system in terms of the observable load voltage $\delta V$, which can be held approximately constant even in the transient regime by connecting a capacitor in parallel to the load, and which we scan over the interval $0,V_{o}]$. Due to the time-dependent nature of the setup, we set $V_{o}$ equal to the mean value of the extracted Seebeck voltages shown in Fig. 8 [87]. This captures both the short- and open-circuit limits, where the efficiency necessarily vanishes due to zero extracted power. The efficiency reaches a maximum at an intermediate value of $\delta V$, allowing a full characterization of the junction performance based solely on measurable quantities.
@@
We observe that the transiently enhanced efficiency can clearly surpass the stationary value. It should be noted, however, that the calculated efficiencies are relatively low ($\lesssim 10$%), which is attributed to the specific molecular structure, the coupling configuration, and the applied temperature gradient. For voltages outside the range $0,V_{o}]$, energy is effectively pumped into the system, and the efficiency drops significantly or becomes negative, as dictated by the direction of currents and the form of Eq. (39). More optimized conditions for efficient thermoelectric energy conversion could be explored within the parameter space of this model, or by considering alternative molecular junction setups, and we defer more thorough investigations of quantum thermodynamics phenomena [108, 109] to future work. Ultimately, the Carnot efficiency $\eta_{C}=1-T_{R}/T_{L}$ sets the fundamental upper limit. This bound is only reached in the idealized limit where the thermoelectric figure of merit $ZT\to\infty$; for any finite $ZT$, the efficiency remains strictly below $\eta_{C}$ [110, 111, 87].
@@ VII Conclusion
Applying the GKBA iteratively in the reconstruction equation resulted in a set of 14 coupled equations of motion, referred to as the iiGKBA scheme. Starting from the full iiGKBA scheme, different levels of approximation were explored. Neglecting correction terms in the collision integral reduced to the GKBA framework, still incorporating retarded/advanced propagators beyond the wide-band limit. We exemplified the different levels of approximation by focusing on thermoelectric transport in a two-terminal quantum dot device, for which analytical benchmark results were available. The conventional GKBA proved insufficient for energy current calculations due to an unphysical divergence in the WBLA, which was resolved by the finite-bandwidth iiGKBA scheme. We also found that the short-time transient behavior of the energy current exhibited an abrupt jump, which was correctly resolved by wider-bandwidth Lorentzians in the iiGKBA scheme.
Finally, we modeled thermoelectric transport in a cyclobutadiene molecular junction — an interacting system — analyzing the non-equilibrium dynamics under both weak and strong coupling, as well as varying temperatures. By comparing with the numerical solution of the full Kadanoff-Baym equations, the iiGKBA scheme was found to systematically improve upon GKBA by more accurately capturing relaxation dynamics and steady-state currents. Using the transient charge and heat currents, we investigated thermoelectric energy conversion by extracting the time-resolved Seebeck coefficient (thermopower) and the device efficiency. The molecular junction, when operating in the transient regime, exhibited a clear enhancement in thermoelectric conversion efficiency compared to its stationary-state performance.